OpenStudy (anonymous):

What is the pattern in these numbers 1, 2, 7, 20, 61, 182

OpenStudy (anonymous):

The next one OS either the previous one times 3 then plus one or minus one

OpenStudy (anonymous):

3 +1 is next pattern is *3-1 *3+1

Directrix (directrix):

0, 1, 2, 7, 20, 61, 182, 547 3(0) + 2(1) = 2; 3(1) + 2(2) = 7; 3(2) + 2(7) = 20; 3(7) + 2(20) = 61; 3(20) + 2(61) = 182; 3(61) + 2(182) = 547; @parmiss The above is not my work and comes from this site: http://www.syvum.com/cgi/online/serve.cgi/iq/ar_series4.html?table Read more about the sequence and its nth term there.